OpenStudy (anonymous):

simplify by factoring. \[\sqrt{90}\] i get 3√10

OpenStudy (cwrw238):

sqrt90 = sqrt9 * sqrt 10 you are correct

OpenStudy (unklerhaukus):

√10=√2√5

OpenStudy (anonymous):

we think of number that go into 90 that give is a 'nice ' square root that we can work out easily with no decimals if possible by 'nice' square roots we mean \[\sqrt{4},\sqrt{9},\sqrt{16},\sqrt{25}.....\] these numbers are called 'square numbers because you can form a square out of 4 dots, 9 dots, 16 dots etc...so \[\sqrt{90}=\sqrt{9}\sqrt{10}=3\sqrt{10}\]